# Bangkok 10330, Thailand

### v1

@@ 4.3 Consistency relation with different horizon exit times
$\mathcal{B}(k_{1},k_{2},k_{3})_{k_{1}\rightarrow 0}=-\mathcal{P}_{k_{1}}^{\sharp}\left(3+k_{2}\frac{\partial}{\partial k_{2}}\right)\mathcal{P}_{k_{2}}^{\flat}\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right],$ (44)
as it was originally guessed Maldacena:2002vr . Here $\mathcal{P}_{k_{1}}^{\sharp}$ and $\mathcal{P}_{k_{1}}^{\flat}$ are (38) where $(H^{*},\varepsilon_{i}^{*},\tau^{*})$ is replaced by $(H^{\sharp},\varepsilon_{i}^{\sharp},\tau^{\sharp})$ and $(H^{\flat},\varepsilon_{i}^{\flat},\tau^{\flat})$, respectively. However, we want to argue that this distinction does not modify the result. Indeed, the power spectrum can be computed explicitly at NLO in terms of the Hubble and slow-roll parameters at $\tau$ as
$\mathcal{P}_{k}=\frac{H^{2}}{4k^{3}\varepsilon_{1}}\left[1+\varepsilon_{2}-(2\varepsilon_{1}+\varepsilon_{2})\left(\gamma_{E}-1+\ln(-2k\tau)\right)\right]+\mathcal{O}\big{(}\varepsilon,\tau\big{)},$ (45)
@@
$\mathcal{P}_{k}=\frac{H^{*2}}{4k^{3}\varepsilon_{1}^{*}}\left[1+\varepsilon_{2}^{*}-(2\varepsilon_{1}^{*}+\varepsilon_{2}^{*})\left(\gamma_{E}-1+\ln(-2k\tau^{*})\right)\right]+\mathcal{O}\big{(}\varepsilon^{*},\tau\big{)},$ (46)
which is indeed (39) up to NLO. In general, the power spectrum is expected to take the following form:
$\mathcal{P}_{k}=\frac{H^{*2}}{4k^{3}\varepsilon_{1}^{*}}\sum\limits_{n=0}^{\infty}\mathcal{A}_{n}\big{(}\ln(-k\tau^{*})\big{)}^{n}+\mathcal{O}\big{(}\varepsilon^{*2},\tau\big{)}$ (47)
with $\mathcal{A}_{n}$ some polynomials of the $\varepsilon_{i}^{*}$, when computed around an arbitrary pivot scale [^8] This was explicitly checked at N3LO in Auclair:2022yxs ; Bianchi:2024qyp ; Antoniadis:2024abm .. Therefore, the dependence on $\tau$, the time at which the 2-point function is actually computed, is subleading in the late-time limit. Consequently (44) can be put in the form
$\mathcal{B}(k_{1},k_{2},k_{3})_{k_{1}\rightarrow 0}=-\mathcal{P}_{k_{1}}^{*}\left(3+k_{2}\frac{\partial}{\partial k_{2}}\right)\mathcal{P}_{k_{2}}^{*}\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}},\tau^{\sharp},\tau^{\flat}\right)\right],$ (48)
@@ 5 Conclusion
We have seen that despite the infrared diverging contributions at large physical distances and in the squeezed limits, the consistency relation (1) relating the squeezed limit of the bispectrum to the power spectrum holds at late-time, at NLO in slow-roll parameters for single field inflation models, where the squeezed limit of the amplitude of the bispectrum is given by (37) with any pivot time $\tau^{*}$ that covers the time $\tau$ of the correlator itself. Note that the consistency relation has been verified even in the presence of the small but finite cutoff $k_{1}$ in the form $\ln(k_{1}/k)$ as demonstrated in (37). In the course of the verification, we have shown that the bispectrum at late time (16) computed in Antoniadis:2024abm is actually time independent due to the cancellation of the time dependence of $H,\varepsilon_{i}$ and the logarithm $\ln(-\tau)$ in (16), and also that it is independent of the pivot time at which $H,\varepsilon_{i}$ are evaluated. We also have shown that the power spectrum also has exactly the same property: it is independent of the time at which the 2pt function is defined and of the choice of the pivot time at which the Hubble and slow-roll parameters are defined.
@@
$\mathcal{B}(k_{1},k_{2},k_{3})=\frac{H^{*4}}{32\varepsilon_{1}^{*2}}\frac{\mathscr{A}^{*}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}$ (49)
is finite in this limit at LO in $\varepsilon^{*}$. However, this is not true at higher orders. Indeed, we have found at NLO two logarithms, $\ln(-k_{\text{t}}\tau)$ that diverges in the large physical distance limit and $\ln(k_{1}/k)$ that diverges in the squeezed limit Antoniadis:2024abm . Therefore, the usual $f_{\text{NL}}$ parameter defined by (28), giving, with $k=k_{2}=k_{3}+\mathcal{O}(k_{1})$,
$\begin{split}f_{\text{NL},k_{1}\rightarrow 0}&=\frac{5}{12}\Bigg{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}+2\varepsilon_{1}^{*2}+5\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\\ &\qquad~{}-\big{(}8\varepsilon_{1}^{*2}+6\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\big{)}\big{[}\gamma_{E}-1+\ln(-2k\tau^{*})\big{]}\\ &\qquad~{}-\big{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}\big{)}^{2}\ln\left(\frac{k_{1}}{k}\right)\Bigg{)}+\mathcal{O}\left(\frac{k_{1}}{k},\varepsilon^{*3},\tau\right),\end{split}$ (50)

### v2

@@ 4.3 Consistency relation with different horizon exit times
$\mathcal{B}(k_{1},k_{2},k_{3})_{k_{1}\rightarrow 0}=-\mathcal{P}_{k_{1}}^{\sharp}\left(3+k_{2}\frac{\partial}{\partial k_{2}}\right)\mathcal{P}_{k_{2}}^{\flat}\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}}\right)\right],$ (4.9)
as it was originally guessed [5]. Here $\mathcal{P}_{k_{1}}^{\sharp}$ and $\mathcal{P}_{k_{1}}^{\flat}$ are (4.3) where $(H^{*},\varepsilon_{i}^{*},\tau^{*})$ is replaced by $(H^{\sharp},\varepsilon_{i}^{\sharp},\tau^{\sharp})$ and $(H^{\flat},\varepsilon_{i}^{\flat},\tau^{\flat})$, respectively. However, we want to argue that this distinction does not modify the result. Indeed, the power spectrum can be computed explicitly at NLO in terms of the Hubble and slow-roll parameters at $\tau$ as
$\mathcal{P}_{k}=\frac{H^{2}}{4k^{3}\varepsilon_{1}}\left[1+\varepsilon_{2}-(2\varepsilon_{1}+\varepsilon_{2})\left(\gamma_{E}-1+\ln(-2k\tau)\right)\right]+\mathcal{O}\big{(}\varepsilon,\tau\big{)},$ (4.10)
@@
$\mathcal{P}_{k}=\frac{H^{*2}}{4k^{3}\varepsilon_{1}^{*}}\left[1+\varepsilon_{2}^{*}-(2\varepsilon_{1}^{*}+\varepsilon_{2}^{*})\left(\gamma_{E}-1+\ln(-2k\tau^{*})\right)\right]+\mathcal{O}\big{(}\varepsilon^{*},\tau\big{)},$ (4.11)
which is indeed (4.4) up to NLO. In general, the power spectrum is expected to take the following form:
$\mathcal{P}_{k}=\frac{H^{*2}}{4k^{3}\varepsilon_{1}^{*}}\sum\limits_{n=0}^{\infty}\mathcal{A}_{n}\big{(}\ln(-k\tau^{*})\big{)}^{n}+\mathcal{O}\big{(}\varepsilon^{*2},\tau\big{)}$ (4.12)
with $\mathcal{A}_{n}$ some polynomials of the $\varepsilon_{i}^{*}$, when computed around an arbitrary pivot scale [^8] This was explicitly checked at N3LO in [19, 20, 4].. Therefore, the dependence on $\tau$, the time at which the 2-point function is actually computed, is subleading in the late-time limit. Consequently (4.9) can be put in the form
$\mathcal{B}(k_{1},k_{2},k_{3})_{k_{1}\rightarrow 0}=-\mathcal{P}_{k_{1}}^{*}\left(3+k_{2}\frac{\partial}{\partial k_{2}}\right)\mathcal{P}_{k_{2}}^{*}\left[1+\mathcal{O}\left(\frac{k_{1}}{k_{2}},\tau^{\sharp},\tau^{\flat}\right)\right],$ (4.13)
@@ 5 Conclusion
We have seen that despite the infrared diverging contributions at large physical distances and in the squeezed limits, the consistency relation (1) relating the squeezed limit of the bispectrum to the power spectrum holds at late-time, at NLO in slow-roll parameters for single field inflation models, where the squeezed limit of the amplitude of the bispectrum is given by (4.2) with any pivot time $\tau^{*}$ that covers the time $\tau$ of the correlator itself. Note that the consistency relation has been verified even in the presence of the small but finite cutoff $k_{1}$ in the form $\ln(k_{1}/k)$ as demonstrated in (4.2). In the course of the verification, we have shown that the bispectrum at late time (3.1) computed in [4] is actually time independent due to the cancellation of the time dependence of $H,\varepsilon_{i}$ and the logarithm $\ln(-\tau)$ in (3.1), and also that it is independent of the pivot time at which $H,\varepsilon_{i}$ are evaluated. We also have shown that the power spectrum also has exactly the same property: it is independent of the time at which the 2pt function is defined and of the choice of the pivot time at which the Hubble and slow-roll parameters are defined.
@@
$\mathcal{B}(k_{1},k_{2},k_{3})=\frac{H^{*4}}{32\varepsilon_{1}^{*2}}\frac{\mathscr{A}^{*}}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}$ (5.1)
is finite in this limit at LO in $\varepsilon^{*}$. However, this is not true at higher orders. Indeed, we have found at NLO two logarithms, $\ln(-k_{\text{t}}\tau)$ that diverges in the large physical distance limit and $\ln(k_{1}/k)$ that diverges in the squeezed limit [4]. Therefore, the usual $f_{\text{NL}}$ parameter defined by (3.13), giving, with $k=k_{2}=k_{3}+\mathcal{O}(k_{1})$,
$\begin{split}f_{\text{NL},k_{1}\rightarrow 0}&=\frac{5}{12}\Bigg{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}+2\varepsilon_{1}^{*2}+5\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\\ &\qquad\qquad-\big{(}8\varepsilon_{1}^{*2}+6\varepsilon_{1}^{*}\varepsilon_{2}^{*}+2\varepsilon_{2}^{*2}-\varepsilon_{2}^{*}\varepsilon_{3}^{*}\big{)}\big{[}\gamma_{E}-1+\ln(-2k\tau^{*})\big{]}\\ &\qquad\qquad-\big{(}2\varepsilon_{1}^{*}+\varepsilon_{2}^{*}\big{)}^{2}\ln\left(\frac{k_{1}}{k}\right)\Bigg{)}+\mathcal{O}\left(\frac{k_{1}}{k},\varepsilon^{*3},\tau\right),\end{split}$ (5.2)